OpenStudy (anonymous):

Solve for x: |x + 2| + 16 = 14 (absolute value equations)

OpenStudy (anonymous):

\[\left| x \right| = 5 \implies x = + 5 or x = -5\]

OpenStudy (callisto):

No solution?! |x + 2| + 16 = 14 Subtract 16 from both sides |x + 2| = -2 => Wrong!!!

OpenStudy (anonymous):

yepp :)

OpenStudy (anonymous):

how is -2 wrong?

OpenStudy (callisto):

Absolute value means the positive value of something.. Let say x=-1 |x| = |-1| =1 So, |x+2| should give a positive value... Now, it's negative, so, it's wrong :| At least no x-int. for it :|

OpenStudy (callisto):

See this: http://www.wolframalpha.com/input/?i=%7Cx+%2B+2%7C+%2B+16+%3D+14

OpenStudy (anonymous):

modulus always yields a positive value. thats how it is defined

OpenStudy (anonymous):

but wouldn't you take the absolute value of -2 and turn it to 2?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

-2 doesnt hv a modulus. why would u put it there??

OpenStudy (callisto):

Not in this question at least You've got |x+2| = -2 As I've said, after you use that sign, you should get a +ve value, now, it's negative. So, it's not true..

OpenStudy (anonymous):

modulus is on x+2 and not on -2

OpenStudy (anonymous):

oh! i see what i did wrong. Thanks

OpenStudy (callisto):

Welcome!~